OpenStudy (dls):

If x=1+y+y^2+....infinity ,y=?

OpenStudy (dls):

@hartnn

OpenStudy (ash2326):

if \[|y|<1\] then \[x=\frac{1}{1-y}\] Can you find y from this in terms of x??

OpenStudy (anonymous):

it is sum to infinite G.P

OpenStudy (dls):

how do I know if y is less than one or greater than that?

OpenStudy (dls):

im feeling embarrassed

OpenStudy (anonymous):

\[x = 1 + y(1+y + y^2...)\]

OpenStudy (dls):

lol

OpenStudy (ash2326):

if \(y\ge 1\) then \[x=\infty\]

OpenStudy (dls):

\[\LARGE 1-y=\frac{1}{x}=>y=1-\frac{1}{x}=>\frac{x-1}{x}?\]

OpenStudy (ash2326):

yeah :)

OpenStudy (dls):

aww I did it wrong in exam D:

OpenStudy (ash2326):

oops :(

OpenStudy (dls):

\[\frac{x}{1-x}\] i did this >.<

OpenStudy (dls):

thanks guys :o